In 5 years, Dad will be three times as old as his daughter Jill will be then. If the sum of their present ages is 50, how old are they now?
If x + 5 is Jill's age five years from now, which of the following equations could be used to solve the problem?

Answer by gonzo(654)

i'm not sure what equations you are looking at, but the following equation will solve the problem.
let D = dad's age now.
let J = jill's age now.
D+5 = dad's age 5 years from now.
J+5 = jill's age 5 years from now.
problem states that the sum of their ages now is 50, so
D+J=50
problem states that in 5 years dad will be 3 times as old as jill will be then, so
(D+5) = 3 * (J+5)
to solve this problem, we remove one of the unknowns by expressing it in terms of the other unknown in one of the equations.
i took D+J=50 because it was simpler. i could have used the other equation as well. either way it will work.
if D+J=50, then D=(50-J)
taking that relationship and substituting for D in the other equation, it becomes
50 - J + 5 = 3 * (J + 5)
this becomes
50 - J + 5 = (3 * J) + (3 * 5)
this becomes
55 - J = (3 * J) + 15
this becomes
55 - 15 = (3 * J) + J
which becomes
40 = 4*J
which becomes
10 = J which is the same as J = 10.
if J = 10, then D = 50 - J becomes D = 50 - 10 becomes D = 40
so jill is 10 years old today and dad is 40 years old today.
5 years from now, jill will be 15 years old and dad will be 45 years old.
45 is 3 * 15 so the requirements of the equation is satisfied.
if you substitute y for D and x for J, then these equations become...
(x+y) = 50
(y+5) = 3*(x+5)
y is dad's age now.
x is jill's age now.